OpenStudy (anonymous):

integrate (-6e^2t)/(1+e^4t) iam having lots of trouble with this problem, any formula i try is wrong, any help that includes some steps would be amazing!!!

OpenStudy (amistre64):

i wonder if long division would help :)

OpenStudy (amistre64):

\[\frac{-6e^{2t}}{1+e^{4t} }\] -6e^2t + 6e^6t - 6e^10t ------------------- 1+e^4t ) -6e^2t (-6e^2t - 6e^6t) -------------- 6e^6t (6e^6t+6e^10) ------------- -6e^10t (-6e^10t -6e^14t) ------------------- .....

OpenStudy (amistre64):

\[\frac{-6e^{2t}}{1+e^{4t} }=\sum_{n=0}^{inf}(-1)^{n+1}6e^{(2+4n)t}\] right?

OpenStudy (anonymous):

if the answer is the last part of ur first box, they answer is incorrect. I also typed it into wolfram alpha and it still didnt take it that answer

OpenStudy (amistre64):

wolf says another way to write it is -3 sech(2t) if you can integrate that?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sum+%28%28-1%29%5E%28n%2B1%29e%5E%28%282%2B4n%29t%29%29+from+0+to+infinity i tool out my 6 and it says this is equal to -1/2 sech (2t) lol

OpenStudy (amistre64):

\[6\int \frac{-e^{2t}}{1+e^{4t} }=\int 6\sum_{n=0}^{inf}(-1)^{n+1}e^{(2+4n)t}\]

OpenStudy (amistre64):

\[\int 6(-e^{2t} +e^{6t} - e^{10t}+e^{14t}...)dt=6(-\frac{1}{2}e^{2t} +\frac{1}{6}e^{6t} -\frac{1}{10} e^{10t}+\frac{1}{14}e^{14t}...)\]

myininaya (myininaya):

\[\text{ Let } u=e^{2t} => du=2 e^{2t} dt => -3 du=-6 e^{2t} dt \] \[\int\limits_{}^{}\frac{-3 du}{1+u^2}=\] this should help

OpenStudy (amistre64):

thats not how orsene would do it :)

myininaya (myininaya):

\[-3\tan^{-1}(u)+C=-3\tan^{-1}(e^{2t})+C\]

OpenStudy (anonymous):

i tried sub but i didnt know how to find it int, once i took the derivative how e^4t become u^2?

myininaya (myininaya):

\[u=e^{2t} => u^2=(e^{2t})^2=e^{2 \cdot 2 \cdot t}=e^{4t}\]

myininaya (myininaya):

By the law of exponents of course :)

myininaya (myininaya):

But you can try amistre's way or whoever that orsene guy is which is probably amistre anyways...

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sum+%28%28-1%29%5E%28n%2B1%29*6e%5E%28%282%2B4n%29t%29%2F%28%282%2B4n%29t%29%29+from+0+to+infinity i got that too lol -3 tan-1(2t) ---------- tho i think +C t

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

i put a spurious little t in the denominator by accident

myininaya (myininaya):

little amistre do you mean e^(2t) inside your glamours tan inverse function?

OpenStudy (amistre64):

yes yes yes ... these fat stubby fingers just werent meant for typing

myininaya (myininaya):

glamorous*

OpenStudy (amistre64):

id give your teacher or whoever is giving you this my answer and see f they accept it lol

OpenStudy (phi):

Are you suppose to use these substitutions: u = e^t and t= ln u so dt= du/u

OpenStudy (anonymous):

i love u guyss thank u!

myininaya (myininaya):

i used u=e^{2t}

myininaya (myininaya):

seems to work just peachy! :)

OpenStudy (amistre64):

i used .... infinite series :)

OpenStudy (phi):

I think you get something that gets you an arctangent

myininaya (myininaya):

amistre was trying to invent a new kind of space travel with some weird math

OpenStudy (phi):

The pied piper, eh?

OpenStudy (amistre64):

zarkon wouldve appreciated it; but jamesj would have shook his head in a contemptible manner and scolded me .... im sure

myininaya (myininaya):

amistre I approve your message. :)